OpenStudy (jade27p):

Write the equation of the line that passes through (2, 4) and has a slope of –1 in point-slope form. @mathmale

OpenStudy (chrishogan):

Still need help?

OpenStudy (jade27p):

yes

OpenStudy (chrishogan):

Well Mathmale is here now. They'll help ya.

OpenStudy (jade27p):

thank you for offering :)

OpenStudy (mathmale):

"Write the equation of the line that passes through (2, 4) and has a slope of –1 in point-slope form." Got to know that "point-slope" form of the eq'n of a straight line: y = mx + b, where m is the slope and b is the y-intercept. From the point (2,4) we get x=2 and y=4. m=-1. Subst. these 3 values into y=mx+b. If you do this correctly, you should find it easy to determine the y-intercept, b. Try it, please.

OpenStudy (jade27p):

x-2=-1

OpenStudy (jade27p):

y becomes -4

OpenStudy (jade27p):

so x-2=-1(y-4)

OpenStudy (triciaal):

what happen to your 2 helpers?

OpenStudy (jade27p):

i dont know

OpenStudy (triciaal):

the slope is -1 slope is the same for any 2 points on that line slope is the change in y over the change in x

OpenStudy (triciaal):